OpenStudy (anonymous):

Find the vertex of the graph of the quadratic formula: f(x)=-2x^2+4x-8

OpenStudy (anonymous):

use the fact that the first coordinate of the vertex is \(-\frac{b}{2a}\) in your case it is \(-\frac{4}{2\times -2}=1\)

OpenStudy (anonymous):

(1,-6)